Summers Corp. currently has an EPS of $2.40, and the benchmark PE for the company is 23. Earnings are expected to grow at 5 pPlease also include work in the BA 2 Plus Calculator if possible with this question.
Answer #1

EPS=$2.40

Benchmark PE for the company =23

Earnings are expected to grow at 5% per year. G=0.05

a. Estimate of the current stock price= Benchmark PE for the company * EPS = 23 * $2.40 = $55.2

b.

Revised EPS = EPS(1+G)= $2.40(1+0.05)=$2.52

Target stock price in one year = Benchmark PE for the company* Revised EPS = 23*2.52 =$57.96

C. Implied Return = (Target stock price in one year-Estimate of the current stock price) / Estimate of the current stock price

=(57.96 - 55.2) /55.2 = 2.76/55.2 = 0.05 or 5%